Find the number of oxygen molecules in 0.1 kg, whose speeds lie between 195 and 205 m/s at




0C.

1
Expert's answer
2022-02-04T07:43:03-0500

Solution;

Given;

T=O°c=273K

m=0.1kg which is contained in 3.125mols

"v_1=195"

"v_2=205"

"v=\frac{v_1+v_2}{2}=\frac{195+205}{2}=200m/s"

The number of molecules is;

"dN=G(v)dvN"

But;

"G(v)dv=4πv^2(\frac{M}{2πRT})^{\frac32}e^{\frac{-Mv^2}{2RT}}dv"

Compute each term;

"4\u03c0v^2=4\u03c0\u00d7200^2=160000\u03c0"

"\frac{M}{2πRT}=\frac{0.032}{2π×8.314×273}=2.244×10^{-6}"

"e^{\frac{-Mv^2}{2RT}}=e^{\frac{-0.032×200^2}{2×8.314×273}}=0.7543"

"DV=205-195=10"

"G(v)dv=160000\u03c0\u00d72.244\u00d710^{-6}\u00d70.7543\u00d710=8.5082"

Then the number of molecules are;

"dN=8.5082\u00d73.125\u00d76.02\u00d710^{23}=1.6\u00d710^{25} molecules"